OpenStudy (anonymous):

what is the exact value of sine 15 degrees and cosine 15 degrees

OpenStudy (anonymous):

\[\sin(15^{o})=\frac{ \sqrt{6}-\sqrt{2} }{ 4 }\]

OpenStudy (anonymous):

\[\cos(15^{0})=\frac{ \sqrt{6}+\sqrt{2} }{ 4 }\]

hero (hero):

@donny471, don't just give the answer. 1. Explain the process 2. Provide general formulas 3. Provide steps whenever possible 4. Include the user in the answering process.

OpenStudy (anonymous):

thanks guys but I would like to know how and why it is the answer. Please explain

OpenStudy (anonymous):

Do you know how to use the unit circle?